OpenStudy (anonymous):

Can someone please explain how to graph 1/2cot(x/2)?

OpenStudy (anonymous):

Only 2 periods

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you allowed to use a graphing calculator?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

ok, you'll have to make a table of x,y pairs so you can plot points

jimthompson5910 (jim_thompson5910):

so when x = pi/2, y is... y = 1/2cot(x/2) y = 1/2cot( (pi/2)/2) y = 1/2*cot(pi/4) y = 1/2*1 y = 1/2 So the point (pi/2, 1/2) is on the graph

jimthompson5910 (jim_thompson5910):

do that for other allowed x values in the domain of cot(x) keep in mind that cot(x) = cos(x)/sin(x)

OpenStudy (anonymous):

also my teacher wants us to state the points we are transforming. what are those?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean by that

jimthompson5910 (jim_thompson5910):

are you given a list of points?

OpenStudy (anonymous):

no. like these are the points for a parent tan function: (-pi/2, undefined) (pi/4, -1) (0,0) (pi/4, 1) (pi/2, undefined)

OpenStudy (anonymous):

I don't know how those were formed. they were just given.

jimthompson5910 (jim_thompson5910):

oh gotcha

jimthompson5910 (jim_thompson5910):

well tangent = sine/cosine

jimthompson5910 (jim_thompson5910):

tangent is undefined when cosine is 0 cosine is 0 when the angle is pi/2 or 3pi/2

OpenStudy (anonymous):

so 1 point would be (0, pi/2)?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

well when x = 0, we have cot(x/2) = cos(x/2)/sin(x/2) cot(0/2) = cos(0/2)/sin(0/2) cot(0) = cos(0)/sin(0) cot(0) = 1/0 cot(0) = undefined

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np